Ted is 12 years younger than Francisco. Francisco is 3 times as old as Ted. How old is fransisco
Answer by ikleyn(52781)

.

    F = 3*(F-12)


    F = 3F - 36


    36 = 3F - F


    36 = 2F


     F = 36/2 = 18.


ANSWER.  Fred is 18 years old.  Ted is 18-12 = 6 years old.

Solved.